# Testing Most Influential Sets

### v2

@@ 6 Conclusion
Theoretical results are elaborated upon in the Appendix, where proofs are elaborated upon via several Lemmata. Datasets can be obtained from the cited sources, and code that generates the results will be made available.
@@
Large language models were used to (i) aid and polish writing, (ii) retrieve and discover related work, and (iii) check results for apparent mistakes.
Appendix A1 Exact Influence Formulas
Proposition 2.
Consider the influence in Equation 3. We can simplify this expression to
$\Delta\left(\{i\}\right)=\frac{x_{i}r_{i}}{\sum_{n\neq i}x_{n}^{2}}.$
Proof.
Let $A_{i}\mathrel{\raisebox{0.0pt}{:}}=x_{i}^{2}$ and $D\coloneqq\sum_{n}^{N}X_{n}^{2},$ such that $D_{-\{i\}}=\sum_{n\neq i}^{N}x_{n}^{2}.$ Then
$\Delta\left(\{i\}\right)=\frac{\frac{\sqrt{A_{i}}r_{i}}{(A_{i}+D_{-\{i\}})}}{1-\frac{A_{i}}{(A_{i}+D_{-\{i\}})}}$ $=\frac{(A_{i}^{1.5}\cdot r_{i}+A_{i}^{0.5}\cdot r_{i}\cdot D_{-\{i\}})}{D_{-\{i\}}\left(A_{i}+D_{-\{i\}}\right)}$
$=\frac{r_{i}\cdot x_{i}\cdot\left(A_{i}+D_{-\{i\}}\right)}{D_{-\{i\}}\cdot\left(A_{i}+D_{-\{i\}}\right)}=\frac{r_{i}\cdot x_{i}}{\sum_{n\neq i}^{N}x_{n}^{2}}.$
A1.1 Recursion
It helps to know the influence on the residual, leverage, and full hat matrix:
$(r_{i})_{-\{j\}}$ $=r_{i}+x_{i}\,\Delta\left(\{j\}\right),$
$(h_{i})_{-\{j\}}$ $=x_{i}^{2}/\sum_{n\neq j}x_{n}^{2},$
$(h_{ij})_{-\{k\}}$ $=h_{ij}+\frac{h_{ik}h_{kj}}{1-h_{k}}.$
Proposition 3.
Let ${\mathbb{S}}=\left\{1,\dots,K\right\}.$ Then we can recursively define $\Delta\left({\mathbb{S}}\right)$ as
$\Delta\left({\mathbb{S}}\right)$ $=\hat{\beta}-\hat{\beta}_{-{\mathbb{S}}}$
$=\Delta\left(\{1\}\right)+\Delta\left(\{2\}\right)_{-\{1\}}+\dots+\Delta\left(\{K\}\right)_{-{\mathbb{S}}\setminus K}.$
Proof.
Notice that
[…]
$=\Delta\left(\left\{j\right\}\right)+\Delta\left(\left\{i\right\}\right)_{-\{j\}}.$
Equivalent results trivially hold for larger sets. ∎
Now, we can provide the full details for the induction step in the proof of Proposition 1:
Proof.
$\Delta\left({\mathbb{S}}\right)=\frac{\sum_{k=1}^{K-1}x_{k}r_{k}}{D_{-{\mathbb{S}}}+x_{k}^{2}}+\frac{x_{k}r_{k}}{D_{-{\mathbb{S}}}}+\frac{x_{k}^{2}\sum_{k=1}^{K-1}x_{k}r_{k}}{D_{-{\mathbb{S}}}\left(D_{-{\mathbb{S}}}+x_{k}^{2}\right)}=\frac{\sum_{k=1}^{K}x_{k}r_{k}}{D_{-{\mathbb{S}}}},$
and
$\Delta\left({\mathbb{S}}\right)$ $=\frac{\sum_{k=1}^{K}x_{k}r_{k}}{D_{-{\mathbb{S}}}+x_{K+1}^{2}}+\frac{x_{K+1}r_{K+1}}{D_{-{\mathbb{S}}}}+\frac{x_{K+1}^{2}\sum_{k=1}^{K}x_{k}r_{k}}{D_{-{\mathbb{S}}}\left(D_{-{\mathbb{S}}}+x_{K+1}^{2}\right)}$
$=\frac{\sum_{i=k}^{K}x_{k}r_{k}\left(D_{-{\mathbb{S}}}+x_{K+1}^{2}\right)}{D_{-{\mathbb{S}}}\left(D_{-{\mathbb{S}}}+x_{K+1}^{2}\right)}+\frac{x_{K+1}r_{K+1}}{D_{-{\mathbb{S}}}}=\frac{\sum_{k=1}^{K+1}x_{k}r_{k}}{D_{-{\mathbb{S}}}}.$
@@
    $\operatorname{Var}(X_{1}^{2})=\sigma^{2}>0$ (non-degenerate variance of squares)
@@
where $\mu=\mathbb{E}[X_{1}^{2}]$ and $\sigma^{2}=\operatorname{Var}(X_{1}^{2})$ (finite by $\mathbb{E}[X_{1}^{4}]<\infty$).
@@
For notational simplicity let $S\coloneqq\sum_{i\in{\mathbb{S}}}X_{i}\cdot R_{i}$ and $T\coloneqq D^{-1}_{-{\mathbb{S}}},$ where $D=\sum_{n}X_{n}^{2}$. It holds for any realization $S=s\in\mathbb{R}$ and $T=t\in\mathbb{R^{+}}$. Further, let $\mathrm{MDA}(H)$ denote the maximum domain of attraction of an EVD $H$ where we write $Z\in\mathrm{MDA}(H)$. We specifically denote the Fréchet as $\Phi_{\alpha}$ and the Gumbel as $\Lambda.$ We are interested in the EVD of $\Delta=S\cdot T.$
A3.1 If $S\in\mathrm{MDA}(\Lambda)$
@@
Let $T\in\mathrm{MDA}(\Lambda)$ and $S\in\mathrm{MDA}(\Phi_{a})$ with tail-coefficient $a>0$ and $S$ and $T$ being independent, then $\Delta\left({\mathbb{S}}\right)=S\cdot T\in\mathrm{MDA}(\Phi_{a})$ .
@@
Recall that for Gumbel tails ($S$) the survival function decays double-exponentially, i.e.,
$\mathbb{P}(S>s)\sim\exp\left(-\exp\left(\frac{s-\mu}{\beta}\right)\right)\quad\text{as }s\to\infty,$
@@
We are interested in the EVD of $\Delta$, i.e., $\mathbb{P}(\Delta>\delta)$. Since $\Delta=S\cdot T$ and $S$, $T$ are independent by construction:
@@
$\int_{0}^{\delta/M}\mathbb{P}(S>u)u^{a-1}L_{T}\left(\frac{\delta}{u}\right)\,\mathrm{du}\sim L_{T}(\delta)\int_{0}^{\infty}\mathbb{P}(S>u)u^{a-1}\,\mathrm{du}.$ (A1)
@@
The term in parentheses is slowly varying in $\delta$ since $L_{T}(\delta)$ is slowly varying which concludes the proof. ∎
To summarize, the survival function $\mathbb{P}(\Delta>\delta)$ is regularly varying with index a,-a, and therefore, $\Delta$ has Fréchet tails with tail-coefficient $a.$
@@
$\Delta\left({\mathbb{S}}\right)\mid X_{-{\mathbb{S}}}\in\mathrm{MDA}(E),$
@@
    Choose $\delta$ small enough such that $k=a-1+\delta>0,$ and
@@
  • For $u\geqslant\delta/M$, we have $\delta/u\leqslant M$ s.t. is bounded on compact sets: $L_{T}(\delta/u)\leqslant C_{M}.$
@@
Appendix A5 Estimation
Figure A1: Visualization of Table 1. Most notable is the regime change between thin and polynomial tails. While for the Frechét cases the average MLE is statistically insignificantly different form another, they are all individually different from the Gumbel case. While the finite sample $D^{-1}_{-{\mathbb{S}}}$ has thicker than Normal tails, acting as a regularizer on $\Delta^{\max}$ and thereby resulting in slight underestimation of the theoretically suggested $\xi^{-1}=0.20$.
Table A1: Inverse Shape Estimates for Different Distributions and Sample Sizes
@@
100 Normal–Normal $0.0416$ $0.0319$ 0.0104-0.0104 $0.0425$ $0.0954$
t(5)t(5)–Normal $0.1432$ $0.0363$ $0.0814$ $0.1431$ $0.2028$
Normal–$t(5)$ $0.1706$ $0.0362$ $0.1110$ $0.1697$ $0.2316$
$t(5)t(5)$ $0.2120$ $0.0384$ $0.1497$ $0.2112$ $0.2758$
200 Normal–Normal $0.0267$ $0.0318$ 0.0249-0.0249 $0.0272$ $0.0786$
t(5)t(5)–Normal $0.1394$ $0.0361$ $0.0774$ $0.1400$ $0.1967$
Normal–$t(5)$ $0.1666$ $0.0353$ $0.1091$ $0.1670$ $0.2241$
$t(5)t(5)$ $0.2033$ $0.0374$ $0.1406$ $0.2049$ $0.2627$
500 Normal–Normal $0.0147$ $0.0316$ 0.0358-0.0358 $0.0159$ $0.0635$
t(5)t(5)–Normal $0.1509$ $0.0387$ $0.0854$ $0.1528$ $0.2133$
Normal–$t(5)$ $0.1692$ $0.0367$ $0.1076$ $0.1696$ $0.2316$
$t(5)t(5)$ $0.2088$ $0.0382$ $0.1437$ $0.2096$ $0.2685$
1000 Normal–Normal $0.0101$ $0.0333$ 0.0451-0.0451 $0.0100$ $0.0641$
t(5)t(5)–Normal $0.1596$ $0.0378$ $0.0941$ $0.1603$ $0.2215$
Normal–$t(5)$ $0.1726$ $0.0360$ $0.1137$ $0.1727$ $0.2333$
$t(5)t(5)$ $0.2101$ $0.0374$ $0.1495$ $0.2123$ $0.2689$
2000 Normal–Normal $0.0106$ $0.0327$ 0.0431-0.0431 $0.0114$ $0.0626$
t(5)t(5)–Normal $0.1731$ $0.0371$ $0.1113$ $0.1740$ $0.2323$
Normal–$t(5)$ $0.1759$ $0.0359$ $0.1185$ $0.1762$ $0.2387$
$t(5)t(5)$ $0.2150$ $0.0375$ $0.1521$ $0.2139$ $0.2763$
All results based on 1000 repetitions.
Figure A2: Convergence of cases to the limiting distribution, visualizing Table A1. Solid lines are estimated means from $1000$ repetitions while the shaded area is $\pm 1SD$. The Normal–Normal case as well as the $t(5)t(5)$ case converge quickly to the theoretical $\xi^{-1}=0$ and $\xi^{-1}=0.2$ respectively. The convergence rate of the mixed cases converge slower yet reasonably fast. In particular, the t(5)t(5)–Normal case is visibly requiring more observations to stabilize $D^{-1}_{\mathbb{S}}$. In general this finding supports the relevance and applicability of our limiting results to finite samples.
Figure A3: Simulation exercise for the performance of the simple MLE based on block maxima, correcting for block size. While the corrected location parameter $\hat{\tilde{a}}$ is close to unbiased, the scale parameter $\hat{b}$ suffers some downward bias using simple block maxima, which is in line with Dombry & Ferreira (2019). However, for practical purposes the block maxima is expected to be fitting reasonably well, as visible in panel (a).
Appendix A6 Auxiliary Results for Case Studies
As mentioned in the main text, Table A2 summarizes the results for testing the preselected set and its subsets for significant influence of the percent of black population on the violent crimes committed per population.
Table A2: Influence of % Black Population on Violent Crimes
@@
2nd Partial after excl. 1st 22 0.0241-0.0241 $0.0051$ $0.0022$ $0.0141$

### v3

@@ 6 Conclusion
Proofs are detailed in the Appendix, datasets are from the cited sources, and code to reproduce results are available at https://github.com/konradld/testingMIS.
@@
Large language models were used to (i) aid and polish writing, (ii) discover and retrieve related work, (iii) check results for apparent mistakes.
@@
    $\mathrm{Var}(X_{1}^{2})=\sigma^{2}>0$ (non-degenerate variance of squares)
@@
where $\mu=\mathbb{E}[X_{1}^{2}]$ and $\sigma^{2}=\mathrm{Var}(X_{1}^{2})$ (finite by $\mathbb{E}[X_{1}^{4}]<\infty$).
@@
For notational simplicity let $S\coloneqq\sum_{i\in{\mathbb{S}}}X_{i}\cdot R_{i}$ and $T\coloneqq D^{-1}_{-{\mathbb{S}}},$ where $D=\sum_{n}X_{n}^{2}$. This holds for any realization $S=s\in\mathbb{R}$ and $T=t\in\mathbb{R}^{+}$. Further, let $\mathrm{MDA}(H)$ denote the maximum domain of attraction of an EVD $H$ where we write $Z\in\mathrm{MDA}(H)$. We specifically denote the Fréchet as $\Phi_{\alpha}$ and the Gumbel as $\Lambda.$ We are interested in the EVD of $\Delta=S\cdot T.$
@@
Let $S\in\mathrm{MDA}(\Lambda)$ and $T\in\mathrm{MDA}(\Phi_{a})$ with tail-coefficient $a>0$ and $S$ and $T$ being independent, then $\Delta\left({\mathbb{S}}\right)=S\cdot T\in\mathrm{MDA}(\Phi_{a}).$
@@
Recall that for Gumbel tails ($S$) the survival function decays faster than any polynomial, i.e.,
$\mathbb{P}(S>s)\sim\exp\left(-\frac{s-\mu}{\beta}\right)\quad\text{as }s\to\infty,$
@@
We are interested in the EVD of $\Delta$, i.e., $\mathbb{P}(\Delta>\delta)$. Since $\Delta=S\cdot T$ and $S$, $T$ are independent by assumption:
@@
$\int_{0}^{\delta/M}\mathbb{P}(S>u)u^{a-1}L_{T}\left(\frac{\delta}{u}\right)\,\mathrm{d}u\sim L_{T}(\delta)\int_{0}^{\infty}\mathbb{P}(S>u)u^{a-1}\,\mathrm{d}u.$ (A2.1)
@@
The term in parentheses is slowly varying in $\delta$ since $L_{T}(\delta)$ is slowly varying.
Thus, the survival function $\mathbb{P}(\Delta>\delta)$ is regularly varying with index a,-a, and therefore, $\Delta$ has Fréchet tails with tail-coefficient $a$, which concludes the proof. ∎
@@
$\Delta\left({\mathbb{S}}\right)\mid X_{-{\mathbb{S}}}\in\mathrm{MDA}(E).$
@@
    Choose $\delta$ such that $k=a-1+\delta>0,$ and
@@
  • For $u\geqslant\delta/M$, we have $\delta/u\leqslant M,$ which is bounded on compact sets: $L_{T}(\delta/u)\leqslant C_{M}.$
@@
Appendix A4 Auxiliary Material
@@
20 Normal–Normal $0.1385$ $0.2541$ 0.0206-0.0206 $0.1293$ $0.2902$
t(5)t(5)–Normal $0.2417$ $0.2865$ $0.0863$ $0.2329$ $0.3994$
Normal–$t(5)$ $0.2529$ $0.2479$ $0.0910$ $0.2505$ $0.4089$
$t(5)t(5)$ $0.3271$ $0.2719$ $0.1601$ $0.3144$ $0.4922$
30 Normal–Normal $0.1021$ $0.1803$ 0.0079-0.0079 $0.1101$ $0.2291$
t(5)t(5)–Normal $0.1989$ $0.1845$ $0.0917$ $0.1994$ $0.3067$
Normal–$t(5)$ $0.2023$ $0.1846$ $0.0817$ $0.2101$ $0.3225$
$t(5)t(5)$ $0.2670$ $0.1912$ $0.1453$ $0.2668$ $0.3961$
50 Normal–Normal $0.0668$ $0.1255$ 0.0153-0.0153 $0.0631$ $0.1466$
t(5)t(5)–Normal $0.1708$ $0.1304$ $0.0890$ $0.1701$ $0.2532$
Normal–$t(5)$ $0.1843$ $0.1286$ $0.1015$ $0.1797$ $0.2704$
$t(5)t(5)$ $0.2307$ $0.1363$ $0.1420$ $0.2305$ $0.3203$
75 Normal–Normal $0.0474$ $0.0960$ 0.0163-0.0163 $0.0485$ $0.1176$
t(5)t(5)–Normal $0.1500$ $0.1006$ $0.0821$ $0.1469$ $0.2173$
Normal–$t(5)$ $0.1735$ $0.1057$ $0.1037$ $0.1749$ $0.2434$
$t(5)t(5)$ $0.2214$ $0.1046$ $0.1561$ $0.2213$ $0.2885$
100 Normal–Normal $0.0422$ $0.0788$ 0.0115-0.0115 $0.0429$ $0.0959$
t(5)t(5)–Normal $0.1389$ $0.0846$ $0.0790$ $0.1369$ $0.1948$
Normal–$t(5)$ $0.1721$ $0.0865$ $0.1183$ $0.1731$ $0.2291$
$t(5)t(5)$ $0.2083$ $0.0876$ $0.1492$ $0.2056$ $0.2639$
200 Normal–Normal $0.0253$ $0.0506$ 0.0075-0.0075 $0.0253$ $0.0582$
t(5)t(5)–Normal $0.1418$ $0.0587$ $0.1020$ $0.1403$ $0.1833$
Normal–$t(5)$ $0.1655$ $0.0576$ $0.1272$ $0.1632$ $0.2013$
$t(5)t(5)$ $0.2072$ $0.0595$ $0.1691$ $0.2054$ $0.2450$
300 Normal–Normal $0.0169$ $0.0406$ 0.0096-0.0096 $0.0180$ $0.0455$
t(5)t(5)–Normal $0.1423$ $0.0461$ $0.1122$ $0.1446$ $0.1731$
Normal–$t(5)$ $0.1705$ $0.0463$ $0.1377$ $0.1695$ $0.2006$
$t(5)t(5)$ $0.2082$ $0.0511$ $0.1757$ $0.2082$ $0.2406$
500 Normal–Normal $0.0157$ $0.0319$ 0.0049-0.0049 $0.0154$ $0.0381$
t(5)t(5)–Normal $0.1488$ $0.0378$ $0.1245$ $0.1499$ $0.1739$
Normal–$t(5)$ $0.1667$ $0.0354$ $0.1419$ $0.1662$ $0.1903$
$t(5)t(5)$ $0.2087$ $0.0385$ $0.1833$ $0.2085$ $0.2323$
1000 Normal–Normal $0.0103$ $0.0220$ 0.0042-0.0042 $0.0109$ $0.0243$
t(5)t(5)–Normal $0.1591$ $0.0287$ $0.1408$ $0.1589$ $0.1771$
Normal–$t(5)$ $0.1734$ $0.0252$ $0.1565$ $0.1743$ $0.1905$
$t(5)t(5)$ $0.2105$ $0.0262$ $0.1938$ $0.2101$ $0.2273$
Table A1: Inverse shape estimates for different distributions and sample sizes. They indicate fast convergence to asymptotic predictions ($0.0$ for the Normal-Normal, and $0.2$ for other cases) — for all settings, the predicted value is contained between the 25th and 75th quantile. Estimates are based on 1,000 repetitions.
Figure A1: Simulation exercise for the performance of the simple MLE based on block maxima, correcting for block size. While the corrected location parameter $\hat{\tilde{a}}$ is close to unbiased, the scale parameter $\hat{b}$ suffers some downward bias using simple block maxima, which is in line with Dombry and Ferreira (2019). However, for practical purposes the block maxima are expected to be fitting reasonably well, as visible in panel (a).
@@
2nd after excl. 1st 22 0.0241-0.0241 $0.0051$ $0.0022$ $0.0141$
Table A2: Influence of % Black Population on Violent Crimes. The table summarizes the results for testing the preselected set and its subsets for significant influence of the percent of black population on the violent crimes committed per population.